OpenStudy (anonymous):

(x^3 -3x^2 - 16x - 12) ÷ (x - 6)

OpenStudy (anonymous):

Do you know how to use Synthetic Division?

OpenStudy (anonymous):

This is a great tutorial! http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (anonymous):

is it x^2 + 6x - 2

OpenStudy (anonymous):

\[= (x ^{3}-3x ^{2}-16x-12)/(x-6)\] \[= ((x+2)(x+1)(x-6))/(x-6)\] when \[x \neq 6\] \[=(x+2)(x+1)\]

OpenStudy (anonymous):

@Tymaryland619 Sorry, it is not x^2+6x-2

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (anonymous):

@dhanushka has the correct answer.

OpenStudy (anonymous):

@Mrbonez321 Thank you :-)

OpenStudy (anonymous):

i neede it in that form you wrote

OpenStudy (anonymous):

x^2+3x+2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@Tymaryland619 Take a look at the Purple Math tutorial for future reference.

OpenStudy (anonymous):

ok i'm logging out